OpenStudy (anonymous):

Can someone please help me with a polynomial function problem? Since my pre-cal final is after this holiday break, my teacher gave us a study packet to work on and turn in once we return. After two days I've finished 199 out of the 200 problems. And i'm stuck on this one problem. A candle mold shaped like a rectangular prism has a square base. Its height is 5 centimeters longer than its width. The mold has a volume of 144 cubic centimeters. How wide is the mold? I don't even know where to start D: Pre-cal is by far my worst subject.

OpenStudy (cwrw238):

let x be the width of the base then the height = x+ 5

OpenStudy (cwrw238):

the volume of the prism = w^2h where h = height and w = width so now you can form an equation and solve for x

OpenStudy (cwrw238):

w = x and h = x + 5 so we have volume = x^2(x + 5) = 144 solve this for x

Directrix (directrix):

@TeacupsAndMigranes Do you know or have you heard of the Rational Root Theorem? I'm asking because you will need to solve x^3 + 5x - 144 = 0 for x.

OpenStudy (anonymous):

I believe so. That's the x=p/q thing, correct? Where it's the ratios of the factors of constants? And... the leading coefficient?

Directrix (directrix):

Correct.

OpenStudy (anonymous):

So use that on x^3 + 5x - 144 = 0 and I'll be able to find X?

OpenStudy (anonymous):

So my constant term is 144 and my leading coefficient is 1?

Directrix (directrix):

Yes. Take the factors of 144 and place them over the factors of 1 and remember that they could be positive or negative.

Directrix (directrix):

Because the leading term is x^3, you can tell that whatever the root is, it has to be fairly small because you are cubing it.

Directrix (directrix):

For example, 72 is a candidate from the RR Theorem, but once you cube 72, you can tell it is too big to go to zero when added to 5*72 and -144.

Directrix (directrix):

Ask a question if you want to.

OpenStudy (anonymous):

The factors I believe are right is 1,2,3,4,6,8,9,12,16,18,24,36,48,72, and 144. (negative and positive.) So I would plug one of those numbers in for X and see if once solved it would equal 0?

Directrix (directrix):

Yes. If polynomial equation has one or more rational roots, then they will come up.

OpenStudy (anonymous):

Okay, great! Thank you so much for the help! :D